Calculus I, Spring 2024 (Thomas’ Calculus Early Transcendentals 13ed), http://www.math.nthu.edu.tw/˜ wangwc/

Brief solutions to selected problems in homework 01

1. Section 8.8: Solutions, common mistakes and corrections:

Figure 1: Solution to Section 8.3, problem 13